In an electrical circuit two resistors of 2 Ω and 4 Ω respectively are connected in series to a 6 V battery. The heat dissipated by the 4 Ω resistor in 5 s will be [NCERT Exemplar Problems]
(a) 5 J
(b) 10 J
(c) 20 J
(d) 30 J

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
nswer: c Explanation: (c) Total resistance of the combination Rs = 2 + 4 = 6 Ω Current, I = \(\frac{V}{R_{s}}=\frac{6}{6}\)= 1 Heat dissipation in 4 Ω resistor, H = I²Rt= 1² × 4 × 5 = 20 J